True/false:

If the sample size is n = 21 and the population standard deviation is unknown, the critical value t needed to construct a 90% confidence level for μ is t = 1.721.
Answer #1

n = 21

Degrees of freedom = df = n - 1 = 21 - 1 = 20

At 90% confidence level the t is ,

= 1 - 90% = 1 - 0.90 = 0.1

/ 2 = 0.1 / 2 = 0.05

t /2,df = t0.05,20 = 1.724    ( using student t table)

false